CAVITY ENHANCED ABSORPTION SPECTROSCOPY IN THE 0.8 μm REGION

\underline{MAN-CHOR CHAN}, S. H. YEUNG, Department of Chemistry, The Chinese University of Hong Kong, Shatin, N. T. Hong Kong.

Since the pioneering work by O’Keefe and Deacon,^{a} cavity ring down spectroscopy has been widely recognized as a sensitive technique for absorption spectroscopy. A variation of the technique, cavity enhanced absorption spectroscopy (CEAS) combined with phase sensitive detection initiated by Peeters et al.^{b} has been adopted in our laboratory using our newly built high resolution Ti:sapphire laser spectrometer and a 1.5 meter cavity ring down cell sealed with high reflection (>99.97%) concave mirrors. Comparing to the typical CEAS using a fast digital oscilloscope, this implementation improves the signal-to-noise ratio by a factor of 2-3 for the transitions of acetylene in the 0.8 μm region. In this paper, the experimental setup and the performance of our system will be discussed.

aA. O’Keefe and D. A. G. Deacon, Rev. Sci. Instr. 59, 2544 (1988).

bR. Peeters, G. Berden, A. Ólafsson, L. J. J. Laarhoven, and G. Meijer, Chem. Phys. Lett. 337, 231 (1991).